OpenStudy (anonymous):

Is the line for the equation x = 4 horizontal or vertical? What is the slope of the line? Select the best answer. The line is horizontal. The slope is 0. The line is vertical. The slope is undefined. The line is horizontal. The slope is undefined. The line is vertical. The slope is 0. Help needed! (:

OpenStudy (ryguy):

vertical the slope is undefined

OpenStudy (anonymous):

Thanks dude!!!

OpenStudy (ryguy):

no problem :)

OpenStudy (anonymous):

:) Are you up for helping with a few more?

OpenStudy (ryguy):

i can try.

OpenStudy (anonymous):

Select the coordinates of two points on the line y = 8. (8, 8) and (8, 0) (8, -8) and (8, 8) (0, 8) and (8, 0) (0, 8) and (8, 8)

OpenStudy (ryguy):

D

OpenStudy (anonymous):

I thought it was but I was unsure.

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

Last one I need help with

OpenStudy (ryguy):

hold on

OpenStudy (ryguy):

im pretty sure its A